1.) Find the definite integrals a.) sdx b.) S dx x2-4x+9 c.) ſ 4arccos xdx d.) Se-3x sin 5xdx

e.) S In(4 + x2)dx f.) s cosst dt sin t g.) St tan 5 2xsec 4 2xdx h.) ſ sin(-4x)cos3xdx
Answer #1


19 I s dx (x-8) + 8 X-8 S 4x +32 lu(x-8) + C. da b I 4[ j dx + est to) = 4x+8 ln(7-6)}tc. a lesz 0907 = SG3121455) tant (2-2) 32 e sin (53) da 3 da da da Cox (57) 86-37 cl cof(12) ſe-32 => 1 = -(*** sin57) + 5] -2737 (10 (152) – (=sin(sa) 32 3 3 / - 3